### v1

@@ C.4 Proof of Theorem˜3
$\|(\Sigma_{v,t}^{*})^{-1}\|_{2}\ =\frac{1}{\lambda_{\min}(\Sigma_{v,t}^{*})}\leq\frac{1}{\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v}}.$ (41)
Substituting into (39) yields
$\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\frac{\delta_{v}}{\lambda_{\min}(\hat{\Sigma}_{v})\,(\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v})}.$
Congruence with $C$ then gives
$\bigl\|C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C\bigr\|_{2}\leq\|C\|_{2}^{2}\,\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\frac{\|C\|_{2}^{2}\,\delta_{v}}{\lambda_{\min}(\hat{\Sigma}_{v})\,(\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v})}.$ (42)
To guarantee $C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C\preceq\phi_{v}I$, it is sufficient that the right-hand side of (42) does not exceed $\phi_{v}$. This requirement is equivalent to $\delta_{v}\leq\delta_{v}^{\max}(\phi_{v})$ with $\delta_{v}^{\max}(\phi_{v})$ defined in (25). Moreover, any $\delta_{v}$ satisfying this bound automatically satisfies the condition $\delta_{v}<\lambda_{\min}(\hat{\Sigma}_{v})$ used in (40) and (41), ensuring consistency of the argument.
$\delta_{v}\leq\theta_{v}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})}\theta_{v}.$
Thus, it is sufficient to require that this upper bound does not exceed $\delta_{v}^{\max}(\phi_{v})$, which guarantees (27). Solving the resulting quadratic inequality in $\theta_{v}$ yields the explicit condition in (26).
$(ii)$ Process uncertainty bound. Define $Y_{t}:=X_{t}^{1/2}\Delta_{t}X_{t}^{1/2}\succeq 0$. Then,
$X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}=X_{t}^{1/2}Y_{t}(I+Y_{t})^{-1}X_{t}^{1/2}.$
and since $Y_{t}\succeq 0$,
$\|Y_{t}(I+Y_{t})^{-1}\|_{2}=\lambda_{\max}(Y_{t})/(1+\lambda_{\max}(Y_{t}))=\|Y_{t}\|_{2}/(1+\|Y_{t}\|_{2}).$
Therefore,
$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\|X_{t}^{1/2}\|_{2}^{2}\cdot\frac{\|Y_{t}\|_{2}}{1+\|Y_{t}\|_{2}}.$
However,
$\|Y_{t}\|_{2}=\|X_{t}^{1/2}\Delta_{t}X_{t}^{1/2}\|_{2}\leq\|X_{t}\|_{2}\|\Delta_{t}\|_{2},$
resulting in the bound
$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\frac{\|X_{t}\|_{2}^{2}\,\|\Delta_{t}\|_{2}}{1+\|X_{t}\|_{2}\|\Delta_{t}\|_{2}}.$ (43)
Now, for the minimum-norm construction of $\Delta_{t}$ in (22),
$\|\Delta_{t}\|_{2}\leq\|A^{+}\|_{2}^{2}\,\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}.$ (44)
By Lemma˜4 in Appendix C.3, $\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}\leq\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}$. On the other hand,
$\begin{split}\|X_{t}\|_{2}&\leq\|(\Sigma_{x,t}^{-})^{-1}\|_{2}+\|C\|_{2}^{2}\|(\Sigma_{v,t}^{*})^{-1}\|_{2}\\ &=\frac{1}{\lambda_{\min}(\Sigma_{x,t}^{-})}+\frac{\|C\|_{2}^{2}}{\lambda_{\min}(\Sigma_{v,t}^{*})}.\end{split}$
By Corollary˜2, $\lambda_{\min}(\Sigma_{v,t}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{v})$. Moreover, since by definition
$\Sigma_{x,t}^{-}=A\Sigma_{x,t-1}A^{\top}+\Sigma_{w,t-1}^{*}\succeq\Sigma_{w,t-1}^{*},$
we have $\lambda_{\min}(\Sigma_{x,t}^{-})\geq\lambda_{\min}(\Sigma_{w,t-1}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{w})$. Thus
$\|X_{t}\|_{2}\leq\delta_{x}^{\max}$ (45)
with $\delta_{x}^{\max}$ as defined in (28).
By the monotonicity of $f(x,y):=x^{2}y/(1+xy)$ on $\real{2}_{+}$, replacing $\|X_{t}\|_{2}$ and $\|\Delta_{t}\|_{2}$ in (43) by their respective upper bounds from (44) and (45) preserves the inequality. We thus obtain
$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\frac{(\delta_{x}^{\max})^{2}\|A^{+}\|_{2}^{2}\,\bigl(\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}\bigr)}{1+\delta_{x}^{\max}\|A^{+}\|_{2}^{2}\,\bigl(\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\,\theta_{w}\bigr)}.$
Requiring the right-hand side to be $\leq\phi_{x}$ is equivalent to
$\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}\leq\frac{\phi_{x}}{\|A^{+}\|_{2}^{2}\,\delta_{x}^{\max}\,(\delta_{x}^{\max}-\phi_{x})},$
which ensures (30). Solving the quadratic inequality yields the explicit bound (29).
Combining $(i)$ and $(ii)$ gives $\Phi_{t}\preceq\phi_{v}I+\phi_{x}I\preceq\phi_{N}I,t\geq q$. Since $\phi_{v}+\phi_{x}\leq\phi_{N}$Proposition˜2 applies: the downsampled DR Riccati map is a contraction in the Thompson part metric. By Banach’s fixed-point theorem, the recursion admits a unique fixed point, and the DR Kalman gain converges.
@@ D.1 Proof of Theorem˜4
The prediction dynamics satisfy
$\tilde{e}_{t}=Ae_{t-1}+w_{t-1}-\hat{w}$.
Eliminating $\tilde{e}_{t}$ yields the posterior error recursion
@@
At steady state, the DRKF covariances satisfy
@@
with
$K_{\infty}=\Sigma_{x,\infty}^{-,*}C^{\top}\!\bigl(C\,\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*}\bigr)^{-1}.$
Using the Joseph form,
$\Sigma_{x,\infty}^{*}\;=\;(I-K_{\infty}C)\,\Sigma_{x,\infty}^{-,*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}.$
@@
$\Sigma_{x,\infty}^{*}$ $=F_{\infty}\,\Sigma_{x,\infty}^{*}\,F_{\infty}^{\top}+(I-K_{\infty}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}.$
Rearranging,
$\Sigma_{x,\infty}^{*}-F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}=(I-K_{\infty}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}\succeq 0.$
Since $\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*}\succ 0$, the right-hand side is in fact $\succ 0$, implying the strict Lyapunov inequality $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}.$ In fact, it implies that $F_{\infty}$ is Schur stable. Indeed, if $F_{\infty}$ had an eigenvalue on or outside the unit circle, no $\Sigma_{x,\infty}^{*}\succ 0$ could satisfy $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}$. [^9] This is the standard equivalence: $F$ is Schur $\Leftrightarrow$ there exists $P\succ 0$ with $P-FPF^{\top}\succ 0$.
With $F_{\infty}$ Schur, the homogeneous error dynamics $e_{t}^{0}=F_{\infty}e_{t-1}^{0}$ thus converge to 0. With noises, the covariance of $e_{t}$ under the least-favorable distributions obeys the steady-state equation derived above and thus converges to $\Sigma_{x,\infty}^{*}$. Taking traces gives the claimed convergence and boundedness.
@@ D.2 Proof of Corollary˜5
Since $F_{\infty}$ is Schur by Theorem˜4, the affine recursion converges, yielding
@@
If $\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$, the constant input term vanishes, and thus $m_{t}\to 0$. ∎
@@ D.3 Proof of Theorem˜5
Proof.
By the strong duality of Wasserstein DRO and the orthogonality principle for quadratic losses, the one-step minimax problem admits a saddle point $(\psi_{t}^{*},\mathbb{P}_{e,t}^{*})$ such that
$\inf_{\psi_{t}\in\mathcal{F}_{t}}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})=\operatorname{Tr}[\Sigma_{x,t}^{*}],$
where $\Sigma_{x,t}^{*}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ is the DRKF posterior covariance.
Thus, for any admissible $\psi_{t}$,
$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})\geq\operatorname{Tr}[\Sigma_{x,t}^{*}],$
almost surely. Since $\Sigma_{x,t}^{*}\to\Sigma_{x,\infty}^{*}$ by Theorem˜3, both lower bounds (31) and (32) follow. For the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}$, the least-favorable posterior covariance is exactly $\Sigma_{x,\infty}^{*}$. Hence, the worst-case one-step MSE equals $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ for all sufficiently large $t$, establishing equality in the pointwise limit and in the long-run average. ∎

### v3

@@ C.4 Proof of Corollary˜3
$\|Z_{t}\|_{2}^{2}\leq\|\hat{\Sigma}_{v}\|_{2}\ \|\Sigma_{v,t}^{*}\|_{2}\leq\|\hat{\Sigma}_{v}\|_{2}\ \bar{\lambda}_{v},$
so $\{Y_{t}\}$ and $\{Z_{t}\}$ are uniformly bounded. Therefore, $\{z_{t}\}$ is bounded, and by the Bolzano–Weierstrass theorem, there exists a subsequence $t_{k}$ such that $z_{t_{k}}\to z_{\infty}$. Writing
$z_{\infty}:=(\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},Y_{\infty},Z_{\infty}),$
we have $(\Sigma_{w,t_{k}-1}^{*},\Sigma_{v,t_{k}}^{*})\to(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ along this subsequence.
Passing to the limit along this subsequence in the constraints of (7), and using $\Sigma_{x,t_{k}-1}\to\Sigma_{x,\infty}$ and the closedness of the PSD cone, shows that $z_{\infty}$ is feasible for the stationary SDP (13).
We now show that $z_{\infty}$ is in fact optimal for (13). Assume $\theta_{w}>0$ and $\theta_{v}>0$. [^16] The degenerate case $\theta_{w}=\theta_{v}=0$ reduces to the nominal KF, for which the statement is immediate. Then (13) satisfies Slater’s condition. For instance, choose $\Sigma_{w,\infty}=\hat{\Sigma}_{w}+\varepsilon I$, $\Sigma_{v,\infty}=\hat{\Sigma}_{v}+\varepsilon I$, $Y=\hat{\Sigma}_{w}$, $Z=\hat{\Sigma}_{v}$, $\Sigma_{x,\infty}=0$, and $\Sigma_{x,\infty}^{-}=\Sigma_{w,\infty}$ for sufficiently small $\varepsilon>0$; this makes all LMIs $\succ 0$ and the trace inequalities strict. Therefore, strong duality holds and the KKT conditions are necessary and sufficient for optimality.
For each $t\geq 1$, let $\lambda_{t}$ be a dual optimal solution associated with the primal optimizer $z_{t}$ of (7) (existence follows from Slater’s theorem). Moreover, since the primal feasible sets are uniformly bounded and Slater’s condition holds, the set of primal–dual optimal pairs is bounded. Thus, after passing to a further subsequence if necessary, we have $\lambda_{t_{k}}\to\lambda_{\infty}$. Taking limits in the KKT conditions of (7) along $t_{k}$ (all primal/dual feasibility constraints are closed, and complementarity is preserved under limits) yields that $(z_{\infty},\lambda_{\infty})$ satisfies the KKT conditions of the stationary SDP (13). Consequently, $z_{\infty}$ is optimal for (13), and $(\Sigma_{x,\infty}^{-},\Sigma_{x,\infty})=(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$.
Finally, by Theorem˜3, the gain sequence $\{K_{t}\}$ converges. Along the subsequence $t_{k}$, $(\Sigma_{x,t_{k}}^{-},\Sigma_{v,t_{k}}^{*})\to(\Sigma_{x,\infty}^{-,*},\Sigma_{v,\infty}^{*})$, and continuity of $K(\Sigma_{x,t}^{-},\Sigma_{v})=\Sigma_{x,t}^{-}C^{\top}(C\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v})^{-1}$ on $\mathbb{S}^{n_{x}}_{++}\times\mathbb{S}^{n_{y}}_{++}$ yields $\lim_{k\to\infty}K_{t_{k}}=K_{\infty}^{*}$. Since $\{K_{t}\}$ has a unique limit, we conclude that $K_{\infty}^{*}=\lim_{t\to\infty}K_{t}$.
@@ D.1 Proof of Theorem˜4
The prediction error evolves as $\tilde{e}_{t}=Ae_{t-1}+w_{t-1}-\hat{w}$. Eliminating $\tilde{e}_{t}$ yields the posterior error recursion
@@
At steady-state, the DRKF covariances satisfy
@@
with $K_{\infty}^{*}=\Sigma_{x,\infty}^{-,*}C^{\top}\!\bigl(C\,\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*}\bigr)^{-1}$. Using the Joseph form,
$\Sigma_{x,\infty}^{*}\;=\;(I-K_{\infty}^{*}C)\,\Sigma_{x,\infty}^{-,*}\,(I-K_{\infty}^{*}C)^{\top}+K_{\infty}^{*}\Sigma_{v,\infty}^{*}(K_{\infty}^{*})^{\top}.$
@@
$\Sigma_{x,\infty}^{*}=F_{\infty}\,\Sigma_{x,\infty}^{*}\,F_{\infty}^{\top}+(I-K_{\infty}^{*}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}^{*}C)^{\top}+K_{\infty}^{*}\Sigma_{v,\infty}^{*}(K_{\infty}^{*})^{\top}.$
Hence,
$\Sigma_{x,\infty}^{*}-F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}=Q_{\infty},$
where $Q_{\infty}:=(I-K_{\infty}^{*}C)\Sigma_{w,\infty}^{*}(I-K_{\infty}^{*}C)^{\top}+K_{\infty}^{*}\Sigma_{v,\infty}^{*}(K_{\infty}^{*})^{\top}$. Since $\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*}\succ 0$, for any $x\neq 0$,
$x^{\top}Q_{\infty}x=\|\Sigma_{w,\infty}^{*,1/2}(I-K_{\infty}^{*}C)^{\top}x\|^{2}+\|\Sigma_{v,\infty}^{*,1/2}(K_{\infty}^{*})^{\top}x\|^{2}>0,$
which implies $Q_{\infty}\succ 0$. Thus, $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}$. By the standard discrete-time Lyapunov characterization, this strict inequality implies that $F_{\infty}$ is Schur stable.
Using the temporal independence of $\{w_{t}\}$ and $\{v_{t}\}$ and their independence from $e_{t-1}$, the error covariance $P_{t}:=\mathrm{cov}(e_{t})$ satisfies
$P_{t}=F_{\infty}P_{t-1}F_{\infty}^{\top}+Q_{\infty},$
whose unique fixed point is $\Sigma_{x,\infty}^{*}$. Since $F_{\infty}$ is Schur, $P_{t}\to\Sigma_{x,\infty}^{*}$ and $\sup_{t\geq 0}\mathbb{E}[\|e_{t}\|^{2}]<\infty$, completing the proof.
@@ D.2 Proof of Corollary˜5
If $\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$, the constant input term vanishes, and thus $m_{t}\to 0$.
@@ D.3 Proof of Theorem˜5
For each $t\geq 0$, define the stage-wise minimax value
$V_{t}\coloneqq\inf_{\psi_{t}\in\mathcal{F}_{t}}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t}),$
where $J_{t}(\psi_{t},\mathbb{P}_{e,t})$ is understood almost surely with respect to $\sigma(\mathcal{Y}_{t-1})$. Under Assumption˜1, Lemma˜1 and Theorem˜1 imply that the minimax problem admits a Gaussian least-favorable distribution and that the corresponding optimal estimator is the DRKF $\psi_{t}^{*}$. Moreover, the minimax value is given by $V_{t}=\operatorname{Tr}[\Sigma_{x,t}]$, where $\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ is the DRKF posterior covariance at time $t$. Consequently, for any admissible estimator $\psi_{t}\in\mathcal{F}_{t}$,
$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})\geq V_{t}=\operatorname{Tr}[\Sigma_{x,t}].$
Taking $\liminf_{t\to\infty}$ and using Theorem˜3, which guarantees $\Sigma_{x,t}\to\Sigma_{x,\infty}^{*}$, yields (27). The corresponding long-run average lower bound (28) follows by Cesàro summability.
We now prove the asymptotic optimality of the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}^{*}$. Let $\bar{x}_{t}^{\infty}:=\psi_{\infty}(\mathcal{Y}_{t})$ and $\bar{x}_{t}^{*}:=\psi_{t}^{*}(\mathcal{Y}_{t})$ denote the posterior estimates produced by the steady-state and time-varying DRKFs, respectively. Define the difference $d_{t}:=\bar{x}_{t}^{\infty}-\bar{x}_{t}^{*}$.
A direct comparison of the measurement updates yields
$d_{t}=F_{\infty}d_{t-1}+\Delta_{t}\eta_{t}^{*},\quad t\geq 1,$ (36)
where $F_{\infty}:=(I-K_{\infty}^{*}C)A$, $\Delta_{t}:=K_{\infty}^{*}-K_{t}$, and $\eta_{t}^{*}:=y_{t}-C{\bar{x}^{-}_{t}}^{*}-\hat{v}$ is the innovation of the time-varying DRKF. By Theorem˜4, $F_{\infty}$ is Schur, and by Theorem˜3 we have $\|\Delta_{t}\|_{2}\to 0$.
[…]
$\begin{split}\mathbb{E}\left[\|\eta_{t}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]\leq 2\|C\|_{2}^{2}\,\mathbb{E}\left[\|x_{t}-{\bar{x}^{-}_{t}}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]+2\,\mathbb{E}\left[\|v_{t}-\hat{v}\|^{2}\mid\mathcal{Y}_{t-1}\right].\end{split}$
Under Assumption˜2, the ambiguity sets have finite Wasserstein radii, so all $\mathbb{P}_{v,t}\in\mathbb{D}_{v,t}$ have uniformly bounded second moments about $\hat{v}$. Moreover, Theorem˜3 implies $\Sigma_{x,t}\to\Sigma_{x,\infty}^{*}$, and hence $\sup_{t\geq 0}\operatorname{Tr}[\Sigma_{x,t}]<\infty$. Since $\Sigma_{x,t}^{-}=A\Sigma_{t-1}A^{\top}+\Sigma_{w,t-1}^{*}$ and $\Sigma_{w,t-1}^{*}\preceq\overline{\lambda}_{w}I$ by Corollary˜2, it follows that $\sup_{t\geq 0}\operatorname{Tr}[\Sigma_{x,t}^{-}]<\infty$, and therefore
$\sup_{t\geq 0}\mathbb{E}[\|x_{t}-{\bar{x}^{-}_{t}}^{*}\|^{2}\mid\mathcal{Y}_{t-1}]<\infty.$
Consequently, there exists $M_{\eta}<\infty$ such that
$\sup_{t\geq 0}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}\mathbb{E}\left[\|\eta_{t}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]\leq M_{\eta}.$
Therefore,
$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}\mathbb{E}\left[\|\Delta_{t}\eta_{t}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]\leq\|\Delta_{t}\|_{2}^{2}\,M_{\eta}\xrightarrow[t\to\infty]{}0.$
Since $F_{\infty}$ is Schur, iterating (36) and using the above uniform bound yields that $d_{t}\to 0$ in mean square; in particular,
$D_{t}:=\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}\mathbb{E}\left[\|d_{t}\|^{2}\mid\mathcal{Y}_{t-1}\right]\xrightarrow[t\to\infty]{}0.$
Let $e_{t}^{*}:=x_{t}-\bar{x}_{t}^{*}$ denote the time-varying DRKF posterior error. Since $x_{t}-\bar{x}_{t}^{\infty}=e_{t}^{*}-d_{t}$, we have
$\|x_{t}-\bar{x}_{t}^{\infty}\|^{2}\leq\|e_{t}^{*}\|^{2}+\|d_{t}\|^{2}+2\|e_{t}^{*}\|\|d_{t}\|.$
Taking conditional expectations and then the supremum over $\mathbb{D}_{e,t}$ yields
$\begin{split}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})\leq\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t}^{*},\mathbb{P}_{e,t})+D_{t}+2\sqrt{D_{t}}\sqrt{\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t}^{*},\mathbb{P}_{e,t})}.\end{split}$
By stage-wise minimax optimality of the time-varying DRKF, $\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t}^{*},\mathbb{P}_{e,t})=V_{t}=\operatorname{Tr}[\Sigma_{x,t}]$. Thus,
$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})\leq V_{t}+D_{t}+2\sqrt{V_{t}}\sqrt{D_{t}}.$ (37)
Taking $\limsup_{t\to\infty}$ and using $D_{t}\to 0$ and $V_{t}\to\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ gives
$\limsup_{t\to\infty}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})\leq\operatorname{Tr}[\Sigma_{x,\infty}^{*}].$
Combined with the lower bound (27) applied to $\psi_{\infty}$, this proves (29). The Cesàro-average optimality statement (30) follows similarly since $\sup_{t}V_{t}<\infty$ and $D_{t}\to 0$ imply $\frac{1}{T}\sum_{t=0}^{T-1}D_{t}\to 0$, and hence all error terms in (37) have vanishing Cesàro mean.